OpenStudy (anonymous):

Solve the following system of equations: 1 / x + 1 / y = 1 / 15 x / y = 5

OpenStudy (ash2326):

We have \[\frac1 x+\frac 1 y=\frac {1}{15}\] and \[\frac{x}{y}=5\] so from second \[\frac{1}{y}=\frac{5}{x}\] Do you get this?

OpenStudy (anonymous):

yes i do, it makes more sense the way you have it written out

OpenStudy (ash2326):

Now we have \[\frac{1}{y}=\frac{5}{x}\] Use this in first \[\frac1 x +\frac 1 y=\frac {1}{15}\] \[\frac1 x +\frac 5 x= \frac 1 {15}\] Now can you find x from this?

OpenStudy (anonymous):

on the left side would it become 6/2x? or just 6/x?

OpenStudy (ash2326):

OK tell me \[\frac 13+ \frac23=?\]

OpenStudy (anonymous):

3/3=1

OpenStudy (anonymous):

okay i got it, and then from there I just solve for x. Thank you

OpenStudy (ash2326):

Yeah, then use x to find y. Let me know if you need help anywhere:)

OpenStudy (anonymous):

I got the right answers! thanks again